OpenStudy (anonymous):

Your realized income is $3,167.89/month, and your fixed expenses are $954.32/every 2 weeks. If you save 50% of your discretionary monies each month, how much are you saving?

OpenStudy (anonymous):

Im taking a test and could use some help for my consumer math class! Homeschooled kid.....

OpenStudy (anonymous):

There are 12 months in a year $43381.52/12 = $3615.12 Max amount per month = $3615.12

OpenStudy (anonymous):

so what do i do for the 50% part? im confused

OpenStudy (amistre64):

isnt RI = DI - Exp therefore DI = RI + exp and you want to know what half of that is right?